Sharon’s Boards sells a snowboard, Xpert, that is popular with snowboard enthusiasts. Information relating to Sharon’s purchases of Xpert snowboards during September is shown below. During the same month, 121 Xpert snowboards were sold. Sharon’s uses a periodic inventory system.

Date

Explanation

Units

Unit Cost

Total Cost

Sept. 1 Inventory

26

$97

$ 2,522

Sept. 12 Purchases

45

105

4,725

Sept. 19 Purchases

20

107

2,140

Sept. 26 Purchases

50

108

5,400

   Totals

141

$14,787

Compute the ending inventory at September 30 and cost of goods sold using the FIFO and LIFO methods.
Answer #1
Ending inventory units 20 =141-121
FIFO:
Ending inventory at September 30 2160 =20*108
Cost of goods sold 12627 =14787-2160
LIFO:
Ending inventory at September 30 1940 =20*97
Cost of goods sold 12847 =14787-1940